# The Impact of Unstated Norms in Bias Analysis of Language Models

### v4

@@ Abstract
Bias in large language models (LLMs) has many forms, from overt discrimination to implicit stereotypes. Counterfactual bias evaluation is a widely used approach to quantifying bias and often relies on template-based probes that explicitly state group membership. It measures whether the outcome of a task performed by an LLM is invariant to a change in group membership. In this work, we find that template-based probes can lead to unrealistic bias measurements. For example, LLMs appear to mistakenly cast text associated with White race as negative at higher rates than other groups. We hypothesize that this arises artificially via a mismatch between commonly unstated norms, in the form of markedness, in the pretraining text of LLMs (e.g., Black president vs. president) and templates used for bias measurement (e.g., Black president vs. White president). The findings highlight the potential misleading impact of varying group membership through explicit mention in counterfactual bias quantification.
@@ 1 Introduction
There has been a surge of interest in and research on bias in machine-learning models. An area of focus is the presence of bias in large language models (LLMs), especially those trained on extensive datasets sourced primarily from the internet. These models have attracted increasing attention due to the influence they are poised to have as they are rapidly integrated into a wide array of applications Gallegos et al. (2023); Wan et al. (2023); Sheng et al. (2021); Liu et al. (2023). Bias in these models manifests in diverse ways, ranging from overtly discriminatory generations to more subtle expressions like stereotypes. In particular, biases toward underprivileged groups, such as racial minorities, have rightfully garnered attention, as they persist across many social contexts. Uncovering these issues represents a crucial step in addressing the potential implications of such biases in downstream applications.
@@
Counterfactual bias evaluation is a common approach in bias quantification that measures invariance, or lack thereof, in the outcomes of a model for a particular task across different groups, holding all else equal De-Arteaga et al. (2019); Czarnowska et al. (2021). A pertinent example is perturbing the race associated with a piece of text from one group (e.g. White) to another (e.g. Black) and checking if a model’s sentiment prediction changes. Although this is a widely used approach in bias quantification, it ignores the fact that LLM training data does not necessarily follow the same structure for different groups.
Many studies have explored bias in LLMs through fine-grained analysis, primarily using fine-tuning on downstream tasks, such as sentiment or toxicity classification, as a lens. These studies employ a diverse set of metrics to detect variations in model behavior Gallegos et al. (2023); Delobelle et al. (2022); Czarnowska et al. (2021); Mökander et al. (2023); Liang et al. (2021); Ribeiro et al. (2020); Levy et al. (2023); Echterhoff et al. (2024). As an example, some LLMs have been shown to demonstrate bias by tending to categorize text related to Muslims as toxic Rae et al. (2021). Standard and Chain-of-Thought (CoT) Wei et al. (2024) prompting has also been used for bias quantification and identification in LLMs Ganguli et al. (2023); Cheng et al. (2023); Kaneko et al. (2024); Tian et al. (2023). While some challenges arise in using prompting in this setting Zayed et al. (2024), it remains a useful tool. As a standard approach in counterfactual bias analysis, these works often use template-based probing datasets, but a quantitative study of potential caveats with such datasets has not been reported.
In this work, we find evidence that explicitly varying group membership through templates and measuring response invariance has flaws in text data, particularly when default group membership is often left implied in natural language. The concept of default group membership in the absence of direct assignment has been extensively studied in linguistics under the broader category of markedness Trubetzkoy (1969); Jakobson (1972); Comrie (1986). In sociological contexts, markedness considers the linguistic differences that arise when referring to default groups compared to others. The concept was first extended to social categories, such as gender and race, in Waugh (1982) wherein it is noted that U.S. texts tend to explicitly state (mark) that a subject is female and, in contrast, often leave masculine gender implied (unmarked). That is, it is more common to use the term “CEO” when an individual is male compared to “female CEO” when they are female. Many subsequent studies have affirmed that markedness extends to race and, in particular, that non-White individuals are often referred to along with their race, while White race membership tends to go unstated Cheryan and Markus (2020); Berkel et al. (2017); Brekhus (2002).
@@
As seen below, such presumption around group membership may produce misleading bias measurements if templates that change group membership through group-associated lexicons are used. In this work, we perform counterfactual bias quantification for ternary sentiment-analysis tasks and empirically observe that LLMs unexpectedly demonstrate “bias” against White-race associated texts similar to traditionally underprivileged groups like African Americans. For example, positive or neutral statements associated with the White group are misinterpreted as negative at higher rates than other groups. These patterns are consistent across several bias probing datasets, LLMs, and classification techniques.
Although further investigation is required, we hypothesize that these irregularities are due to the prevalence of markedness in LLM pretraining text rather than a true bias present in the LLMs themselves. A key assumption underlying unmarked representations is that humans are adept at recognizing unstated implications in text. LLMs trained solely on unstructured next-token prediction, which underpins almost all modern LLM pretraining, may lack the ability to perceive such implications, resulting in surprising behavior. Using templates that represent group membership through explicit description likely makes certain text appear uncommon for traditionally unmarked groups. As such, these templates may lead to artificially elevated error rates in LLMs, skewing bias measurements in unpredictable ways.
Overall, the results in this work provide strong quantitative evidence that the common approach of using templates that rely on marking to establish group membership for bias quantification in LLMs has flaws. These flaws make such measurements less reliable lenses into actual bias dynamics. The experiments establish that these issues affect a wide variety of LLMs and are likely intrinsic to the models themselves rather than the particular pipeline being applied.
@@ 2 Related Work
In Blodgett et al. (2021), a critical study of several bias datasets (StereoSet, CrowS-Pairs, WinoBias, WinoGender) revealed systematic issues compromising, to varying degrees, the precision or clarity of biases or stereotyping tendencies of LLMs measured by these datasets. Among other issues, including poor definitions, misalignment, and logical failures, the authors identify out-of-domain text due to markedness as potentially clouding the proposed measurements. The investigation therein is comprehensive and bolsters our hypothesis that markedness plays a significant role in the results below. However, the study does not quantify the effect of these flaws, nor does it explore their impact on downstream task probes, as done here.
@@
Several studies considering the extent to which markedness is incorporated into LLMs (or their multimodal variants) or affects their predictions exist Bender et al. (2021); Wolfe and Caliskan (2022a, b). Each of these studies notes that markedness plays a critical role in the way models make predictions and that these models have internalized aspects of markedness through their training. These studies reveal certain biases related to markedness but do not investigate counterfactual bias or the interaction of markedness with template-based probes.
Reporting bias shares significant conceptual overlap with markedness. It refers to the linguistic phenomenon in which descriptive elements are highlighted only when they deviate from the norm Grice (1975). Generally, it relates to the broader theme of commonsense reasoning in language where statements may not explicitly encompass all the underlying information. This can be problematic for LLMs, which rely on concepts learned from the text. In Shwartz and Choi (2020), it is shown that pretrained LMs are impacted by reporting bias with respect to their estimation of rare events. This work is concerned with reporting bias in so far as it intersects with markedness and has been shown to impact the predictions of LMs.
@@ 3 Methodology
In natural language processing, bias measurement commonly considers disparities with respect to sensitive attributes such as gender or race Czarnowska et al. (2021). Within each sensitive attribute are various protected groups. The groups considered here for the race attribute are American Indian, Asian, African American, Hispanic, Pacific Islander, and White. A standard approach to measuring bias is to evaluate model performance disparities when protected groups are varied. Ideally, model performance is invariant to the change of protected group.
It should be noted that race and ethnicity have specific and distinct anthropological definitions. However, many works and bias datasets use the terms interchangeably or mix the concepts, including those leveraged in this study. For example, the templates in Czarnowska et al. (2021), discussed below, consider the group “Hispanic” under the sensitive attribute of race, but it is commonly considered an ethnicity Lopez et al. (2023). Thus, the term race is used throughout for consistency with previous work despite it constituting an imperfect fit with some of the protected groups.
Below, counterfactual bias quantification is applied to a collection of LLMs. Two sets of experiments are designed. In the first, LLMs are fine-tuned for three-way sentiment classification using the SST5 dataset Socher et al. (2013). Bias is then measured by varying group membership through the use of several template-based datasets. In the second, LLMs are prompted to perform classification directly on the template-based datasets without fine-tuning. For this study, the sensitive attribute of race is considered and disparities in classification of text associated with different racial groups are measured. In both pipelines, analysis is conducted by measuring the false positive rate (FPR) between groups. Three different template-based dataset are used in this work. [^1] All experimental code is found at: https://github.com/VectorInstitute/UnstatedNormsAndLLMBias
@@ 3.1.1 Amazon Dataset
This dataset consists of templates for generating examples for a specific sensitive attribute, such as gender and race, as well as generic templates that may be used to produce examples for any sensitive attribute Czarnowska et al. (2021). In this paper, both templates specific to the attribute of race and generic templates are used. All templates have a sentiment label and are filled with different race-associated adjectives to generate samples coupled to a specific group. Examples are as follows.
@@ 3.1.2 NS-Prompts Dataset
The NS-Prompts dataset is created from templates using sentences sourced from Wikipedia consisting of professions or job titles and their description Alnegheimish et al. (2022). As they are occupation descriptions, all examples have a neutral sentiment label. The original templates have potential formats
@@
where X is a profession or job title and Y is the associated description. A completed example might be “A cartoonist is a person who specializes in drawing cartoons or comics."
@@
“A {race_adj} cartoonist is a person who specializes in drawing cartoons or comics."
While all samples have a neutral label, we are interested in measuring model prediction errors while varying these group adjectives.
@@ 3.1.3 Regard Dataset
In its original form, the dataset of Sheng et al. (2019) is used to measure bias toward specific demographics, including race. The templates consist of phrases joined by “and.” The first phrase, referred to as regard, measures language polarity and social perceptions of a demographic, while the second part, focused solely on sentiment, measures overall language polarity. A positive, neutral, or negative label is applied to each phrase separately. An original template highlighting the regard and sentiment phrases in red and blue, respectively, is as follows.
@@
Labels: Regard: Negative. Sentiment: Positive.
In this work, only the first part of the templates, corresponding to the regard component, is used, along with the regard label. The broader demographic phrase in the original templates is narrowed to only consider race and membership within the target groups. As an example, the template above becomes:
@@ 3.2 Bias Quantification Metric
Negative-Sentiment FPR indicates the percentage of sentences labeled positive or neutral that are incorrectly classified as negative. An elevated Negative-Sentiment FPR gap indicates a potential lack of preference for a group, where such sentences are cast as negative at a higher rate. Alternatively, Positive-Sentiment FPR denotes the frequency with which negative or neutral sentences are incorrectly identified as positive. A Positive-Sentiment FPR gap greater than zero suggests a preference for a group, wherein negative or neutral sentences are construed as positive at a higher rate. In particular, an elevated Negative-Sentiment FPR gap coupled with a Positive-Sentiment FPR gap below zero implies that examples for a group are classified as negative and neutral at a higher rate compared with other groups, suggesting that the group is an unfavorable one for the LLM.

### v5

# Template-Based Probes Are Imperfect Lenses for Counterfactual Bias Evaluation in LLMs

@@ Abstract
Bias in large language models (LLMs) has many forms, from overt discrimination to implicit stereotypes. Counterfactual bias evaluation is a widely used approach to quantifying bias and often relies on template-based probes that explicitly state group membership. It aims to measure whether the outcome of a task performed by an LLM is invariant to a change in group membership. In this work, we find that template-based probes can introduce systematic distortions in bias measurements. Specifically, we consistently find that such probes suggest that LLMs classify text associated with White race as negative at disproportionately elevated rates. This is observed consistently across a large collection of LLMs, over several diverse template-based probes, and with different classification approaches. We hypothesize that this arises artificially due to linguistic asymmetries present in LLM pretraining data, in the form of markedness, (e.g., Black president vs. president) and templates used for bias measurement (e.g., Black president vs. White president). These findings highlight the need for more rigorous methodologies in counterfactual bias evaluation, ensuring that observed disparities reflect genuine biases rather than artifacts of linguistic conventions.
@@ 1 Introduction
There has been a surge of interest in, and research on, bias in machine learning models. An important area of focus is the presence of bias in large language models (LLMs), especially those trained on extensive datasets sourced primarily from the internet. These models have attracted increasing attention due to their rapid integration into a wide array of applications (Gallegos et al., 2024; Wan et al., 2023; Sheng et al., 2021; Liu et al., 2023). Bias in these models manifests in diverse ways, ranging from overtly discriminatory generations to more subtle expressions like perpetuating stereotypes. In particular, biases toward underprivileged groups, such as racial minorities, have rightfully garnered attention, as they persist across many social contexts. Uncovering these issues represents a crucial step in addressing the potential implications of such biases in downstream applications.
@@
Counterfactual bias evaluation is a common approach in bias quantification that measures invariance, or lack thereof, in the outcomes of a model for a particular task across different groups, holding all else equal (De-Arteaga et al., 2019; Czarnowska et al., 2021; Martinková et al., 2023; Cimitan et al., 2024). A pertinent example is perturbing the race associated with a piece of text from one group (e.g. White) to another (e.g. Black) and measuring whether a model’s sentiment prediction changes. Although this is a widely used approach in bias quantification, it ignores the fact that LLM training data does not necessarily follow the same structure for different groups.
@@
In this work, counterfactual bias quantification experiments are performed spanning several ternary sentiment-analysis tasks. A wide range of LLMs is considered, and two classification techniques, fine-tuning and prompting, are applied to perform the classification tasks. Empirically, we observe clear abnormalities such that LLMs assign disproportionately negative sentiment to texts explicitly associated with White race, with similarities to traditionally underprivileged groups like African Americans. For example, positive or neutral statements associated with the White group are misinterpreted as negative at higher rates than other groups. These patterns are consistent across bias probing datasets, LLMs, and classification techniques. Overall, the results demonstrate that template-based bias quantification relying on marking has flaws. These limitations reduce the reliability of such measurements as indicators of actual bias dynamics.
The contributions of this work are summarized as follows.
  • We find evidence that counterfactual bias evaluation using template-based probes introduces systematic distortions in bias measurement. The extent to which template-based probes exhibit measurement flaws is systematically quantified through a wide range of experiments. These distortions undermine the usefulness of such datasets as a lens for bias evaluation.
  • This paper constructs two new template-based probing datasets from existing work to validate the findings across different domains. These datasets, and the associated techniques for their construction, may be used in future experiments.
  • This work provides a strong conjecture as to the underlying cause of the aberrant bias measurements. We hypothesize that these disparities are due to the prevalence of markedness in LLM pretraining text, suggesting new research directions.
@@ 2 Related Work
Many studies have explored bias in LLMs through fine-grained analysis, primarily using fine-tuning on downstream tasks, such as sentiment or toxicity classification, as a lens. These studies employ a diverse set of metrics to detect variations in model behavior (Gallegos et al., 2024; Delobelle et al., 2022; Czarnowska et al., 2021; Mökander et al., 2023; Liang et al., 2021; Ribeiro et al., 2020; Levy et al., 2023; Echterhoff et al., 2024; Rae et al., 2022). Standard and Chain-of-Thought (CoT) (Wei et al., 2024) prompting have also been used for bias quantification and identification in LLMs (Ganguli et al., 2023; Cheng et al., 2023; Kaneko et al., 2024; Tian et al., 2023). While some challenges arise in using prompting in this setting (Zayed et al., 2024), it remains a useful tool. A multitude of studies, including those cited above, use template-based probing datasets to perform counterfactual, extrinsic bias analysis in LLMs (Dixon et al., 2018; Huang et al., 2020; Liang et al., 2021; Blodgett et al., 2021; Delobelle et al., 2022; Martinková et al., 2023; Cimitan et al., 2024). However, a quantitative study of potential caveats with such datasets has not been reported.
@@
In Blodgett et al. (2021), a critical study of several bias datasets (StereoSet, CrowS-Pairs, WinoBias, WinoGender) identified systematic issues likely compromising the precision of biases or stereotyping tendencies of LLMs measured by these datasets. Among other issues, including poor definitions, misalignment, and logical failures, the authors suggest out-of-domain text due to markedness as potentially clouding the proposed measurements. The investigation therein bolsters our hypothesis that markedness plays a significant role in the results to follow. However, their study does not quantify the effect of these flaws. Rather, it simply identifies qualities that may be problematic. Their work also focuses on different datasets than those studied here. Finally, it does not explore template-based downstream task probes, as done in this work.
Flaws associated with intrinsic bias metrics, which aim to identify bias in model representations rather than downstream tasks, have been examined in Delobelle et al. (2022). Their work demonstrates that such measures are not well correlated with extrinsic (downstream) bias measures and even fail to provide consistent results between intrinsic measures. The authors identify poorly designed templates, among other factors, as contributing to the issues with intrinsic bias metrics. However, the results do not consider or quantify issues with template-based probes for extrinsic bias metrics. Moreover, their work focuses exclusively on masked language models (MLMs), whereas the experiments below consider both LLMs and MLMs.
2.1 Linguistic Markedness
The concept of default group membership in the absence of direct assignment has been extensively studied in linguistics under the category of markedness (Trubetzkoy, 1969; Jakobson, 1972; Comrie, 1986). In sociological contexts, markedness considers the linguistic differences that arise when referring to default groups compared to others. The idea was first extended to social categories, such as gender and race, in Waugh (1982) wherein it is noted that U.S. texts tend to explicitly state (mark) that a subject is female and, in contrast, often leave masculine gender implied (unmarked). That is, it is more common to use the term “CEO” when an individual is male compared to “female CEO” when they are female. Many subsequent studies have affirmed that markedness extends to race and, in particular, that non-White individuals are often referred to along with their race, while White-race membership tends to go unstated (Cheryan and Markus, 2020; Berkel et al., 2017; Brekhus, 2002).
@@ 2.1 Linguistic Markedness
Several studies considering the extent to which markedness or reporting bias are incorporated into LLMs or affects their predictions exist (Bender et al., 2021; Wolfe and Caliskan, 2022a; b; Cheng et al., 2023; Shwartz and Choi, 2020). Each of these studies notes that markedness plays a critical role in the way models make predictions and that these models have internalized aspects of markedness through their training. These studies reveal certain biases related to markedness or reporting bias but do not investigate counterfactual bias or template-based probes from this perspective. In Section 5, we conjecture that the irregularities observed in the results to follow are driven by markedness in LLM pretraining data.
@@ 3 Methodology
In natural language processing, bias measurement typically examines disparities across sensitive attributes such as gender or race (Czarnowska et al., 2021). Each attribute is composed of various protected groups. Herein, the attribute of race is specifically considered. Within the sensitive attribute of race, we restrict our focus to the protected groups of American Indian, Asian, African American, Hispanic, Pacific Islander, and White. A standard bias measurement approach evaluates model performance disparities when protected groups are varied. Ideally, model performance remains invariant to group changes or substitutions.
It should be noted that race and ethnicity have distinct anthropological definitions, yet many studies and bias datasets use the terms interchangeably, including those used in the experiments to follow. For instance, the templates in Czarnowska et al. (2021), discussed below, categorize “Hispanic” under race, though it is commonly considered an ethnicity (Lopez et al., 2023). To maintain consistency with prior work, the term “race” is used throughout, despite its imperfect fit for some protected groups.
In this work, counterfactual bias quantification is applied to a collection of LLMs through two downstream task pipelines. In the first, LLMs are fine-tuned for three-way sentiment classification using the SST5 dataset (Socher et al., 2013), and bias is measured by varying group membership across multiple template-based datasets. In the second, LLMs classify template-based datasets directly, without fine-tuning, through prompting. As this study primarily examines race as the sensitive attribute, we measure classification performance disparities across racial groups. Both pipelines analyze false positive rate (FPR) discrepancies between groups. Three template-based datasets are used and detailed in the sections to follow. [^1] All experimental code is found at: https://github.com/VectorInstitute/UnstatedNormsAndLLMBias
@@ 3.1 Template-Based Datasets
The construction of the three template-based probing datasets used in the experiments is described in this section. Additional details around the composition of the datasets, including the resulting label distributions, are found in Appendix E.
@@ 3.1.1 Amazon Dataset
This dataset consists of templates for generating examples for a specific sensitive attribute, such as gender and race, as well as generic templates that may be used to produce examples for any sensitive attribute (Czarnowska et al., 2021). Templates specific to the attribute of race and generic templates are both used for the experiments. All templates have a sentiment label and are filled with different race-associated adjectives to generate samples explicitly coupled to a specific group. Examples are as follows.
@@ 3.1.2 NS-Prompts Dataset
The NS-Prompts dataset is created from templates using sentences sourced from Wikipedia consisting of professions or job titles and their description (Alnegheimish et al., 2022). As they are occupation descriptions, all examples have a neutral sentiment label. The original templates have the following potential formats.
@@
where X is a profession or job title and Y is the associated description. A completed example might be “A cartoonist is a person who specializes in drawing cartoons or comics.”
@@
“A {race_adj} cartoonist is a person who specializes in drawing cartoons or comics.”
Though all samples have a neutral label, we are interested in measuring model prediction errors while varying these group adjectives.
@@ 3.1.3 Regard Dataset
The dataset of Sheng et al. (2019) measures bias toward specific demographics, including race. Its templates consist of complete phrases joined by “and.” The first phrase, termed regard, assesses language polarity and social sentiment of a demographic included in the sentence, while the second, originally referred to as sentiment, measures overall language polarity. Each phrase receives a positive, neutral, or negative sentiment label. An original template highlighting regard and sentiment phrases in red and blue, respectively, is
@@
Regard Label: Negative. Sentiment Label: Positive.
Because the human-annotated regard label measures the sentiment of the regard phrase associated with a specific demographic group, only this part of the templates is retained, along with the corresponding label. Going forward, these labels are simply referred to as sentiment. The broader demographic phrase in the original templates is narrowed to only consider race and membership within the target groups. As an example, the template above becomes:
@@ 3.2 Bias Quantification Metric
Negative-Sentiment FPR measures the percentage of positive or neutral sentences misclassified as negative. An elevated Negative-Sentiment FPR gap suggests a potential lack of preference for a group, where such sentences are classified as negative more often. Conversely, Positive-Sentiment FPR denotes the rate at which negative or neutral sentences are misclassified as positive. A Positive-Sentiment FPR gap above zero suggests a preference for a group, where negative or neutral sentences are classified as positive more frequently. An elevated Negative-Sentiment FPR gap combined with a Positive-Sentiment FPR gap below zero indicates that a group’s examples are classified as negative or neutral more often than others, suggesting the group is viewed unfavorably by the LLM.
When considering Negative- or Positive-Sentiment FPR, the interpretation of model errors is fairly straightforward, as discussed above. On the other hand, Neutral-Sentiment FPR is more convolved. Such errors correspond to neutral predictions for samples with either negative or positive labels. Errors are a mix of predictions viewing samples with negative labels more positively and those viewing samples with positive labels more negatively. This clouds interpretation of Neutral-Sentiment FPR gaps without further decomposition of the metric. As such, results are limited to Negative- and Positive-Sentiment FPR in this work.
Note that when using FPR, the metric defined in Equation 1 is a derivative of False Positive Equality Difference (FPED), a standard bias metric (Dixon et al., 2018; Czarnowska et al., 2021). FPED is defined as $\sum|\text{FPR}(X)-\text{FPR}|$, where the sum is over all protected groups and FPR represents the FPR for all samples combined. To allow for more granular representation of performance differences across protected groups, three modifications to the FPED metric are present. First, differences are not summed across groups to retain group-specific differences. Second, the directionality of the difference is maintained by shedding the absolute value. Finally, because the number of samples for each protected group is not necessarily equal, mean FPR across groups is computed rather than the all-sample FPR.